### v1

@@ 4.2 Least Squares Quasi-conformal Energy
Apart from Eq.5, the Beltrami equation can be rewritten as
(6) $\begin{bmatrix}v_{y}\\ -v_{x}\end{bmatrix}=\frac{1}{1-|\mu|^{2}}\begin{bmatrix}(\rho-1)^{2}+\tau^{2}&-2\tau\\ -2\tau&(1+\rho)^{2}+\tau^{2}\end{bmatrix}\begin{bmatrix}u_{x}\\ u_{y}\end{bmatrix}$.
Here we assume $|\mu|<1$ and for convenience, we denote the square matrix on the right-hand side by $\mathbf{A}$. Inspired by the least square conformal energy[levy2002least], Qiu [qiu2019computing] proposed the concept least square quasi-conformal (LSQC) energy of a map.
@@
Let $\mu=\rho+ir$ be a complex-valued function defined on the domain $\Omega$. The least squares quasi-conformal energy of the map $z=(x,y)\mapsto(u,v)$ against the Beltrami coefficient $\mu$ is defined to be
(7) $E_{LSQC}(u,v,\mu)=\frac{1}{2}\int_{\Omega}\|P\nabla u+JP\nabla v\|^{2}\,dx\,dy,$
where
$P=\frac{1}{\sqrt{1-|\mu|^{2}}}\begin{bmatrix}1-\rho&-r\\ -r&1+\rho\end{bmatrix},J=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$,
so that $P^{T}P=\mathbf{A}$ as in Eq.6 and $P^{T}JP=J$.
The LSQC energy admits the decomposition
(8) $E_{LSQC}(u,v,\mu)=E_{\mathbf{A}}(u;\Omega)+E_{\mathbf{A}}(u;\Omega)-\mathcal{A}(u,v)$
where
$E_{\mathbf{A}}(u;\Omega)=\frac{1}{2}\int_{\Omega}\nabla u^{T}\mathbf{A}\nabla udxdy=\frac{1}{2}\int_{\Omega}\|P\nabla u\|^{2}dxdy$
is the Dirichlet-type energies of second-order equation $\nabla\cdot(\mathbf{A}\nabla u(z))=0$ derived by the decoupling method [lui2014teichmuller] and
(9) $\mathcal{A}(u,v)=\int_{\Omega}(u_{y}v_{x}-u_{x}v_{y})dxdy.$
is the area of the image.
@@
where $d_{T}=(x_{1}y_{2}-y_{1}x_{2})+(x_{2}y_{3}-y_{2}x_{3})+(x_{3}y_{1}-y_{3}x_{1})$ is twice the area of the triangle. The Beltrami equation (Eq.2) can be rewritten as $(1-\mu)f_{x}+i(1+\mu)f_{y}=0$ and collected as a complex-valued linear system.
@@
$E_{LSQC}(\textbf{U}=(U_{1},\ldots,U_{|\mathcal{V}|}))$ $=\sum\limits_{T\in\mathcal{T}}d_{T}\Big|\left((1-\mu)f_{x}+i(1+\mu)f_{y}\right)\Bigg|_{T}\Big|^{2}$
@@
We decomposed $\mathbf{U}=(\mathbf{U}_{f}^{\top},\mathbf{U}_{p}^{\top})^{\top}$, where $\mathbf{U}_{f}$ are free points, i.e. variables of the optimization problem, and $\mathbf{U}_{f}$ are the points pinned. Similarly, we can decompose $\mathcal{M}$ in blocked matrices as
@@ 5.1 Spherical Beltrami Differential
Before stating a similar theorem on the relation between $M(\mathcal{R})$ and the space of quasiconformal maps between two Riemann surfaces, we give two lemmas.
@@
By Thm.4.1, there is a quasiconformal homeomorphism $f=f^{\alpha}$ defined on each $z_{\alpha}(U_{\alpha})$ such that $f_{\bar{z}}(z)=\mu_{\alpha}(z)f_{z}(z)$. On each open set $U_{\alpha}$ of the given covering of $\mathcal{R}$, we introduce the new chart $w_{\alpha}=f^{\alpha}\circ z_{\alpha}$.
@@
by Eq.10, consider $f^{\alpha}$ and $f^{\beta}\circ(z_{\beta}\circ z_{\alpha}^{-1})$ on the overlapping regions in the $z_{\alpha}(U_{\alpha})$, both are the solutions of the same Beltrami equations (with coefficients $\mu_{\alpha}$. By Thm.4.1, there exists a conformal map $\varphi$ such that $f_{\alpha}=\varphi\circ(f^{\beta}\circ(z_{\beta}\circ z_{\alpha}^{-1}))$, and thus $f_{\alpha\beta}=\varphi$ is conformal.
@@
Let $\mathcal{R}$ and $\mathcal{R^{\prime}}$ be two conformal sturctures of the unit sphere $\mathbb{S}^{2}$, then for any $\mu\in M(\mathcal{R})$, there exists a quasiconformal mapping $f:\mathcal{R}\rightarrow\mathcal{R^{\prime}}$ satisfying the Beltrami equation, which is unique up to the postcomposition of conformal mappings.
@@
Given $\mu\in M(\mathcal{R})$, by lemma 5.3, there exists a new chart $\mathcal{R}^{\mu}$ such that $id:\mathcal{R}\rightarrow\mathcal{R}^{\mu}$ is quasiconformal w.r.t $\mu$. By uniformization theorem, there exists a conformal mapping $f:\mathcal{R}^{\mu}\rightarrow\mathcal{\mathcal{R}^{\prime}}$ and this conformal mapping is the desired quasiconformal mapping from $\mathcal{R}$ to $\mathcal{R}$
@@
By Thm.5.5, we establish a one-to-one correspondence between spherical self-quasiconformal maps and spherical Beltrami differentials, modulo the conformal automorphism group.
@@
Our aim is to achieve spherical self-diffeomorphism optimization with explicit control in the local geometric distortion. Considering the correspondence between spherical Beltrami differentials and spherical self-quasiconformal maps, we may formulate the problem as optimization over these geometric quantities. Through stereographic projection, the diffeomorphic optimization problem for a spherical surface $\mathcal{M}=\{(U_{N},P_{N}),(U_{S},P_{S})\}$ (Eq.1) can be reformulated as the optimization over spherical Beltrami differentials:
@@
This new formulation includes two energy terms: $E_{1}$ drives task-specific objectives while $E_{2}$ regularizes BCs $\mu$ by constraining its magnitude (e.g., minimizing $|\mu|^{2}$) and smoothness (e.g., minimizing $|\nabla\mu|^{2}$). This dual-energy approach enables fine-tuned control over the quasiconformal mapping’s geometric properties without compromising its task-oriented performance.
Due to Cor.5.9, above formulation is equivalent to
@@
where $\mu(f\big|_{U})$ is the Beltrami coefficient induced by the resitrction of $f$ on $U$. The Beltrami differential depends only on the domain surface’s parameterization and for simplicity, in application we express the sphere self-mapping $f$ in Eq.1 as $f\colon\mathbb{S}^{2}=(P_{S}^{-1}(\overline{\mathbb{D}}),P_{S})\cup(P_{N}^{-1}(\overline{\mathbb{D}}),P_{N})\rightarrow\mathbb{S}^{2}=(\Omega_{1},P_{N})\cup(\Omega_{2},P_{S})$, where $P_{S},P_{N}$ are south,north pole stereographic projections and $\Omega_{j}$ are simply connected domains in $\mathbb{R}^{2}$. From this point of view, we can decompose $f$ as the combination of its restriction on the upper and lower hemispheres, respectively. Equivalently, it can be viewed as the combination of two homeomorphisms on different hemispheres, and their images of the equator coincide. Therefore, the formulation is simplified to
@@ 5.2 Spectral Beltrami Network
A key challenge is developing a differentiable tool to compute quasiconformal mappings from Beltrami differentials, enabling gradient-based optimization. Meanwhile, the mapping should be not subject to any Dirichlet type boundary constraints because the shape of the target domain $f^{\mu_{\alpha}}\circ z_{\alpha}(U_{\alpha})$ is usually not prescribed. An important reason why LSQC is preferred in our work is that it implies that geometry of target domain is intrinsically determined by Beltrami coeffcients and unlike decoupling methods [lui2014teichmuller] imposing a Dirichlet type boundary condition. This characteristic aligns seamlessly with the previously outlined Beltrami differential optimization formulation 13 and provides the flexibility to adjust boundary configurations dynamically through the optimization of the Beltrami differential.
Following the notation in [xu2025neural], we can view the numerical LSQC algorithm as an operator $\mathcal{F}$. Let $(V,F)$ be a triangulation of the unit disk, $\mu=\{\mu_{T}:T\in F\}$ be piecewise constant beltrami coefficients, $(p_{i},q_{i})_{i=1,2}$ be the two pairs of selected points in the mesh and their corresponding prescribed destinations in $\mathbb{R}^{2}$, then the numerical solution $f=\mathcal{F}(\mu,p_{1},p_{2},q_{1},q_{2})$ such that $f(p_{i})=q_{i}$ and $(\frac{\partial f}{\partial\bar{z}}/\frac{\partial f}{\partial z})\Big|_{T}\approx\mu_{T}$. For the aim of applying LSQC to gradient-based optimization problems, Spectral Beltrami Network (SBN) is proposed in the pervious paper. Given Beltrami coefficients $\{\mu_{v}:v\in V^{1}\}$ defined for vertices on a prescribed mesh of $\mathbb{D}$ (or equivalently a directed graph) $\mathcal{G}^{1}=(V^{1},E^{1},F^{1})$ and coordinates of two pinned points $(p_{1},p_{2})$, SBN would predict the mapping $f(V^{1})$ such that $f(V^{1})\approx\mathcal{F}(\mu,p_{1},p_{2},p_{1},p_{2})$ where $\mu=\{\mu_{T}:\mu_{T}=\frac{1}{3}\sum_{v\in T}\mu_{v},T\in F^{1}\}$. For capturing short-range and long-range information exchange between distinct vertices in the mesh, SBN couples mesh spectral layers with the multiscale message passing mechanism. The mesh spectral layer is to project the latent feature vectors into a span of low-frequency mesh Laplacian eigenvectors, do the multinode linear transformation in the frequency domain followed by an inverse projection back to the spatial domain, through which global shape context is injected to the latent space. In addition, the multiscale mesh hierarchy can widen the receptive field without exploding memory, improving the efficiency of information exchange through downsampling and upsampling. Fig.1 shows the architecture of SBN. For more details about the network, please refer to [xu2025neural].
@@ 6 Beltrami Optimization On Spherical Topology (BOOST)
Since we have the Spectral Beltrami Network (SBN) which can predict free-boundary diffeomorphisms on the unit disk $\mathbb{D}$, in the following, we introduce the optimization framework BOOST for the general spherical parameterization problem, which incorporates SBN for freely deforming the 2D parameterization domain.
The SBN model serves as a tool for spherical surface parameterization, transforming a hemisphere into a submanifold of the unit sphere. By optimizing the Beltrami coefficients (BCs) and the coordinates of two fixed points—followed by post-composition with scaling, rotation, and translation—we derive optimal mappings tailored to specific tasks.
@@
The optimization needs to ensure: (1) no overlaps when gluing the hemispheres, and (2) minimal boundary mismatch. Specifically, there must be no overlaps around the boundaries of two transformed hemispheres when gluing back, meanwhile the distance between $P_{S}^{-1}(f^{S}(v))$ and $P_{N}^{-1}(f^{N}(v))$ for each $v\in\partial V^{1}$ should be small enough. For the second goal, we propose a boundary matching loss
$\mathcal{L}_{\textbf{bm}}(f^{S},f^{N})=\frac{1}{|\partial V^{1}|}\sum\limits_{v\in V^{1}}\|P_{S}^{-1}(f^{S}(v))-P_{N}^{-1}(f^{N}(v))\|_{2}^{2}$
@@
$\mathcal{L}_{\textbf{folding}}(\widetilde{\mathcal{M}})=\int_{\Omega^{-}}\det(J_{f})dxdy\approx\frac{1}{|F^{1}|}\sum\limits_{T\in F^{1}}\phi(-\det(J_{f}(T)))$
where $\phi$ is the ReLU function, $\Omega^{-}$ is the set of points whose Jacobian determinant is negative and $J_{f}$ is the Jacobian determinant of the mapping $f\colon V^{1}\rightarrow\widetilde{\mathcal{M}}$.
[…]
$\mathcal{L}_{\textbf{bs}}=\sum\limits_{v\in P_{N}\circ P_{S}^{-1}(f^{S}(\partial V^{1}))}\|\Delta^{2}v\|_{2}^{2}+0.1\cdot\|\Delta v\|_{2}^{2}$
where $\Delta v_{i}=\sum_{[i,j]}\omega_{ij}(v_{j}-v_{i})$ and $\omega_{ij}=\frac{1}{2}(\cot\alpha_{ij}+\cot\beta_{ij})$, $\alpha_{ij}$ and $\beta_{ij}$ are the angles opposite to the edge $[i,j]$, and $\Delta^{2}$ is the standard scalar Laplacian applied twice. The double Laplacian term distributes the derivative change smoothly into the interior, preventing tension discontinuities. An illustration of $\mathcal{L}_{\textbf{bm}},\mathcal{L}_{\textbf{folding}}$ and $\mathcal{L}_{\textbf{bs}}$ is given in Fig.2
Figure 2: Illustration of three issues addressed by the loss functions $\mathcal{L}_{\textbf{bm}},\mathcal{L}_{\textbf{folding}}$ and $\mathcal{L}_{\textbf{bs}}$
In all, let us denote the post-composition of similarity transformation by $g(x,\phi,\varphi,r)=\varphi e^{i\phi}x+r$ and after training of SBN is done, we freeze its parameter and obtain $\mathcal{F}_{\theta^{*}}$, whose output $\mathcal{F}_{\theta^{*}}(\{\mu_{v}\},p_{1},p_{2})\approx\mathcal{F}(\{\mu_{F}\},p_{1},p_{2},p_{1},p_{2})$. And we can reformulate the framework 13 and obtain BOOST as follows
@@
where $\theta_{N}=\{\tilde{\mu}_{N},\tilde{p}_{N,1},\tilde{p}_{N,2},\phi_{N},s_{N},r_{N}\}$ and $\theta_{S}=\{\tilde{\mu}_{S},\tilde{p}_{S,1},\tilde{p}_{S,2},\phi_{S},\varphi_{S},r_{S}\}$ are the parameter sets and
@@
The details discussed are summarized in Alg.1.
@@
    optimization parameters $\eta$: $T_{\text{BC}},T_{\text{pin}}$, $\theta_{N}=\{\tilde{\mu}_{N},\tilde{p}_{N,1},\tilde{p}_{N,2},\phi_{N},s_{N},r_{N}\}$ and $\theta_{S}=\{\tilde{\mu}_{S},\tilde{p}_{S,1},\tilde{p}_{S,2},\phi_{S},\varphi_{S},r_{S}\}$,
@@
13:   Compute $\frac{\partial\mathcal{L}_{total}}{\partial\eta}$ for all parameters $\eta$ to be optimized and then update all $\eta$ with the optimizer $\Psi$.

### v2

@@ 4.2 Least Squares Quasi-conformal Energy
We define the unweighted LSQC energy as the $L^{2}$-norm of the Beltrami equation mismatch.
@@
Let $\mu=\rho+\sqrt{-1}\tau$ be a complex-valued function defined on the domain $\Omega$. The LSQC energy of the map $z=(x,y)\mapsto(u,v)$ against the BC $\mu$ is defined to be
$\tilde{E}_{LSQC}(f,\mu)=\int_{\Omega}|f_{\bar{z}}-\mu f_{z}|^{2}dA.$
Equivalently,
(6) $\tilde{E}_{LSQC}(u,v,\mu)=\frac{1}{2}\int_{\Omega}\|\tilde{P}\nabla u+J\tilde{P}\nabla v\|^{2}\,dx\,dy,$
where $\tilde{P}=\begin{bmatrix}1-\rho&-\tau\\ -\tau&1+\rho\end{bmatrix}$.
@@
where $d_{T}=(x_{1}y_{2}-y_{1}x_{2})+(x_{2}y_{3}-y_{2}x_{3})+(x_{3}y_{1}-y_{3}x_{1})$ is twice the area of the triangle. The Beltrami equation (Equation 2) can be rewritten as $(1-\mu)f_{x}+\sqrt{-1}(1+\mu)f_{y}=0$ and collected as a complex-valued linear system.
@@
$E_{LSQC}(\textbf{U}=(U_{1},\ldots,U_{|\mathcal{V}|}))$ $=\sum\limits_{T\in\mathcal{T}}d_{T}\Big|\left((1-\mu)f_{x}+\sqrt{-1}(1+\mu)f_{y}\right)\Bigg|_{T}\Big|^{2}$
@@
We decomposed $\mathbf{U}=(\mathbf{U}_{f}^{\top},\mathbf{U}_{p}^{\top})^{\top}$, where $\mathbf{U}_{f}$ are free points, i.e. variables of the optimization problem, and $\mathbf{U}_{f}$ are the points pinned. Similarly, we can decompose $\mathcal{M}$ in blocked matrices
@@ 5.1 Spherical Beltrami Differential
In this subsection, we introduce the concept of spherical Beltrami diffferential associated to a spherical quasiconformal map. Before establishing the relation between spherical Beltrami differentials and spherical quasiconformal maps, we give two lemmas.
@@
By Theorem 4.1, there is a quasiconformal homeomorphism $f=f^{\alpha}$ defined on each $z_{\alpha}(U_{\alpha})$ such that $f_{\bar{z}}(z)=\mu_{\alpha}(z)f_{z}(z)$. On each open set $U_{\alpha}$ of the given covering of $\mathcal{R}$, we introduce the new chart $w_{\alpha}=f^{\alpha}\circ z_{\alpha}$.
@@
by Equation 7, consider $f^{\alpha}$ and $f^{\beta}\circ(z_{\beta}\circ z_{\alpha}^{-1})$ on the overlapping regions in the $z_{\alpha}(U_{\alpha})$, both are the solutions of the same Beltrami equations (with coefficients $\mu_{\alpha}$. By Theorem 4.1, there exists a conformal map $\varphi$ such that $f_{\alpha}=\varphi\circ(f^{\beta}\circ(z_{\beta}\circ z_{\alpha}^{-1}))$, and thus $f_{\alpha\beta}=\varphi$ is conformal.
@@
Let $\mathcal{R}$ and $\mathcal{R^{\prime}}$ be two conformal structures of the unit sphere $\mathbb{S}^{2}$, then for any $\mu\in M(\mathcal{R})$, there exists a quasiconformal mapping $f:\mathcal{R}\rightarrow\mathcal{R^{\prime}}$ satisfying the Beltrami equation, which is unique up to the postcomposition of conformal mappings.
@@
Given $\mu\in M(\mathcal{R})$, by Lemma 5.3, there exists a new chart $\mathcal{R}^{\mu}$ such that $id:\mathcal{R}\rightarrow\mathcal{R}^{\mu}$ is quasiconformal w.r.t $\mu$. By uniformization theorem, there exists a conformal mapping $f:\mathcal{R}^{\mu}\rightarrow\mathcal{\mathcal{R}^{\prime}}$ and it is the desired quasiconformal mapping from $\mathcal{R}$ to $\mathcal{R}$
@@
Using Theorem 5.5, we can establish a one-to-one correspondence between spherical self-quasiconformal maps and spherical Beltrami differentials, modulo the conformal automorphism group.
@@
Our aim is to achieve spherical self-homeomorphism optimization with explicit control on the local geometric distortion. Considering the correspondence between spherical Beltrami differentials and spherical self-quasiconformal maps, we may formulate the problem as optimization over these geometric quantities. Through stereographic projection, the optimization problem for a spherical surface $\mathcal{M}=\{(U_{N},P_{N}),(U_{S},P_{S})\}$ (Equation 1) can be reformulated as the optimization over spherical Beltrami differentials:
@@
This new formulation includes two energy terms: $E_{1}$ drives task-specific objectives while $E_{2}$ regularizes BCs $\mu$ by constraining its magnitude (e.g. minimizing $|\mu|^{2}$) and smoothness (e.g. minimizing $|\nabla\mu|^{2}$).
Due to Corollary 5.9, the above formulation is equivalent to
@@
where $\mu(f\big|_{U})$ is the Beltrami coefficient induced by the restriction of $f$ on $U$. The Beltrami differential depends only on the parameter domain of the surface parameterization and for simplicity, in application we express the sphere self-mapping $f$ in Equation 1 as $f\colon\mathbb{S}^{2}=(P_{S}^{-1}(\overline{\mathbb{D}}),P_{S})\cup(P_{N}^{-1}(\overline{\mathbb{D}}),P_{N})\rightarrow\mathbb{S}^{2}=(\Omega_{1},P_{N})\cup(\Omega_{2},P_{S})$, where $P_{S},P_{N}$ are south, north pole stereographic projections and $\Omega_{j}$ are simply connected domains in $\mathbb{R}^{2}$. From this point of view, we can decompose $f$ as the combination of its restriction on the upper and lower hemispheres, respectively. Equivalently, it can be viewed as the combination of two homeomorphisms on different hemispheres, and their images of the equator coincide. Therefore, the formulation is simplified to
@@ 5.2 Spectral Beltrami Network
A key challenge is developing a differentiable tool to compute quasiconformal mappings from Beltrami differentials, enabling gradient-based optimization. Meanwhile, the mapping should be not subject to any boundary constraints because the shape of the target domain $f^{\mu_{\alpha}}\circ z_{\alpha}(U_{\alpha})$ is usually not prescribed. An important reason why LSQC is preferred in our work is that it implies that geometry of target domain is intrinsically determined by Beltrami coefficients. This characteristic aligns with the previously outlined Beltrami differential optimization formulation 10 and provides the flexibility to adjust boundary configurations dynamically through the optimization of the Beltrami differential.
Following the notation in [xu2025neural], we can view the numerical LSQC algorithm as an operator $\mathcal{F}$. Let $(V,F)$ be a triangulation of the unit disk, $\mu=\{\mu_{T}:T\in F\}$ be piecewise constant Beltrami coefficients, $(p_{i},q_{i})_{i=1,2}$ be the two pairs of selected points in the mesh and their corresponding prescribed destinations in $\mathbb{R}^{2}$, then the numerical solution $f=\mathcal{F}(\mu,p_{1},p_{2},q_{1},q_{2})$ such that $f(p_{i})=q_{i}$ and $(\frac{\partial f}{\partial\bar{z}}/\frac{\partial f}{\partial z})\Big|_{T}\approx\mu_{T}$. For the aim of applying LSQC to gradient-based optimization problems, Spectral Beltrami Network (SBN) is proposed in the previous work[xu2025neural]. Given Beltrami coefficients $\{\mu_{v}:v\in V^{1}\}$ defined for vertices on a prescribed mesh of $\mathbb{D}$ (or equivalently a directed graph) $\mathcal{G}^{1}=(V^{1},E^{1},F^{1})$ and coordinates of two pinned points $(p_{1},p_{2})$, SBN predicts the mapping $f(V^{1})$ such that $f(V^{1})\approx\mathcal{F}(\mu,p_{1},p_{2},p_{1},p_{2})$ where $\mu=\{\mu_{T}:\mu_{T}=\frac{1}{3}\sum_{v\in T}\mu_{v},T\in F^{1}\}$. Arbitrary target positions $q_{1},q_{2}$ can be recovered by post-composition with a similarity transformation $g$ satisfying $g(p_{j})=q_{j}$ $(j=1,2)$. By Proposition 4.6, $\mathcal{F}(\mu,p_{1},p_{2},q_{1},q_{2})=g\circ\mathcal{F}(\mu,p_{1},p_{2},p_{1},p_{2})$, ensuring no loss of generality. For capturing short-range and long-range information exchange between distinct vertices in the mesh, SBN couples mesh spectral layers with the multiscale message passing mechanism. The mesh spectral layer is to project the latent feature vectors into a span of low-frequency mesh Laplacian eigenvectors, and do the multi-node linear transformation in the frequency domain followed by an inverse projection back to the spatial domain, through which global shape context is injected to the latent space. In addition, the multiscale mesh hierarchy can widen the receptive field without exploding memory, improving the efficiency of information exchange through downsampling and upsampling. Figure 1 shows the architecture of SBN. For more details about the network, please refer to [xu2025neural]. The key contribution of SBN is that 1. through it, the Beltrami field becomes learnable and 2. with the stop-gradient operator, the choice of pinned points is no longer a discrete manner but also differentiable.
@@ 6 Beltrami Optimization On Spherical Topology (BOOST)
The Spectral Beltrami Network (SBN) can predict free-boundary homeomorphisms on the unit disk $\mathbb{D}$. In the following, we introduce the optimization framework BOOST for the general spherical parameterization problem, which incorporates SBN for freely deforming the 2D parameterization domain.
After the training of SBN, we freeze its parameters. The well-trained SBN model serves as a tool for spherical surface parameterization, transforming a hemisphere into a submanifold of the unit sphere. A significant advantage of the SBN is that the pinned points $P=\{p_{1},p_{2}\}$ are treated as trainable parameters rather than fixed constraints, using the stop gradient operator. By optimizing the Beltrami coefficients (BCs) and the coordinates of two fixed points, followed by post-composition with scaling, rotation, and translation, we derive optimal mappings associated to specific tasks.
@@
The optimization needs to ensure: (1) no overlaps when gluing the hemispheres, and (2) minimal boundary mismatch. Specifically, there must be no overlaps around the boundaries of two transformed hemispheres when gluing back, meanwhile the distance between $P_{S}^{-1}(f^{S}(v))$ and $P_{N}^{-1}(f^{N}(v))$ for each $v\in\partial V^{1}$ should be small enough, ideally zero. For the second goal, we propose a boundary matching loss
$\mathcal{L}_{\text{bm}}(f^{S},f^{N})=\frac{1}{|\partial V^{1}|}\sum\limits_{v\in V^{1}}\|P_{S}^{-1}(f^{S}(v))-P_{N}^{-1}(f^{N}(v))\|_{2}^{2}$
To achieve our first aim of guaranteeing non-overlapping spherical parameterizations when gluing two deformed hemispheres, we employ a folding detection and penalization strategy based on consistent face orientation. Given the initial non-overlapping spherical parameterization $\mathbf{x}$, we first establish a consistent orientation for all triangular faces in the mesh. For each face $T$ with vertices indexed as $\{i,j,k\}$, we compute the unnormalized face normal vector $\mathbf{n}_{T}$ via the cross product of edge vector: $\mathbf{n}_{T}=({\mathbf{x}}_{i}-{\mathbf{x}}_{k})\times({\mathbf{x}}_{j}-{\mathbf{x}}_{i})$. The orientation is correct if $\mathbf{n}_{T}\cdot\mathbf{x}_{i}>0$, indicating that the normal vector points outward from the sphere’s center. For faces where this condition is violated, we swap the vertex indices $j$ and $k$ to reverse the winding order, thereby ensuring all faces maintain a consistent outward-pointing orientation. With consistent orientation established, any subsequent spherical mapping can be validated for self-intersections or folding by computing the signed area of each triangular face. The signed area $A_{T}^{s}$ for an oriented face $T=(i,j,k)$ is computed as: $A_{T}^{s}=\frac{1}{2}\|\mathbf{n}_{T}\|_{2}\cdot\text{sign}(\mathbf{n}_{T}\cdot\mathbf{x}_{i})$. Under the corrected orientation, faces maintaining positive signed areas ($A_{T}^{s}>0$) iff they preserve the non-overlapping property. To explicitly discourage folding during optimization, we incorporate a folding penalty term into our loss function:
$\mathcal{L}_{\text{folding}}=\frac{1}{|F^{1}|}\sum_{T\in F^{1}}\max(0,-A_{T}^{s})$
The folding penalty $\mathcal{L}_{\text{folding}}$ is assigned a higher weight than other regularization terms, because bijectivity is the fundamental topological constraint rather than a soft geometric preference. Similar folding penalty terms have been employed in various spherical mapping works as auxiliary losses to discourage self-intersections such as [ren2024sugar]. Notably, our two-hemisphere gluing strategy has a significant advantage in achieving $\mathcal{L}_{\text{folding}}=0$ compared to existing approaches. Since each hemisphere is predicted by SBN which always inherently produces homeomorphic mappings, face flipping often occurs exclusively at the deformed equator where the two charts are overlapped. The majority of faces maintain correct orientation by construction. In contrast, others must regularize all faces across the entire sphere, as inversions can occur anywhere during optimization. Therefore, it is easier to remain the bijectivity of the result mapping in our works than others and empirically final mappings by BOOST have no foldings.
@@
$\mathcal{L}_{\text{bs}}=\sum\limits_{v\in P_{N}\circ P_{S}^{-1}(f^{S}(\partial V^{1}))}\|\Delta^{2}v\|_{2}^{2}+0.1\cdot\|\Delta v\|_{2}^{2}$
where $\Delta v_{i}=\sum_{[i,j]}\omega_{ij}(v_{j}-v_{i})$ and $\omega_{ij}=\frac{1}{2}(\cot\alpha_{ij}+\cot\beta_{ij})$, $\alpha_{ij}$ and $\beta_{ij}$ are the angles opposite to the edge $[i,j]$, and $\Delta^{2}$ is the standard scalar Laplacian applied twice. The double Laplacian term distributes the derivative change smoothly into the interior, preventing tension discontinuities. An illustration of $\mathcal{L}_{\text{bm}},\mathcal{L}_{\text{folding}}$ and $\mathcal{L}_{\text{bs}}$ is given in Figure 2.
Figure 2: Illustration of three issues addressed by the loss functions $\mathcal{L}_{\text{bm}},\mathcal{L}_{\text{folding}}$ and $\mathcal{L}_{\text{bs}}$
In all, let us denote the post-composition of similarity transformation by $g(x,\phi,\varphi,r)=\varphi e^{i\phi}x+r$ and after training of SBN is done, we freeze its parameters and obtain $\mathcal{F}_{\theta^{*}}$, whose output $\mathcal{F}_{\theta^{*}}(\{\mu_{v}\},p_{1},p_{2})\approx\mathcal{F}(\{\mu_{F}\},p_{1},p_{2},p_{1},p_{2})$. And we can reformulate the framework 10 and obtain BOOST as follows
@@
where $\theta_{N}=\{\tilde{\mu}_{N},T_{\text{BC},N},\tilde{p}_{N,1},\tilde{p}_{N,2},T_{\text{pin},N},\phi_{N},s_{N},r_{N}\}$ and $\theta_{S}=\{\tilde{\mu}_{S},T_{\text{BC},S},\tilde{p}_{S,1},\tilde{p}_{S,2},T_{\text{pin},S},$ $\phi_{S},\varphi_{S},r_{S}\}$ are the parameter sets and
@@
The details discussed are summarized in Algorithm 1.
@@
    optimization parameters $\eta$: $\theta_{N}=\{\tilde{\mu}_{N},T_{\text{BC},N},\tilde{p}_{N,1},\tilde{p}_{N,2},T_{\text{pin},N},\phi_{N},s_{N},r_{N}\}$ and $\theta_{S}=\{\tilde{\mu}_{S},T_{\text{BC},S},\tilde{p}_{S,1},\tilde{p}_{S,2},T_{\text{pin},S},$ $\phi_{S},\varphi_{S},r_{S}\}$,
@@
13:   Compute $\nabla_{\eta}\mathcal{L}_{total}$ and update all $\eta$ with the optimizer $\Psi$.